OpenStudy (anonymous):

@perl Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5 h. Car B traveled 5 mph slower than Car A. How fast did Car B travel?

OpenStudy (perl):

that worked

OpenStudy (anonymous):

thanks

OpenStudy (perl):

we can use distance = rate * time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

donused

OpenStudy (anonymous):

4.5 hr x what?

OpenStudy (perl):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5 h. Car B traveled 5 mph slower than Car A. How fast did Car B travel? distance_A = (rate_A) * 4 hours distance_B = (rate_B) * 4.5 hours rate_B = rate_A - 5 distance_A = distance_B

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

is it 55mph?

OpenStudy (perl):

distance_A = (rate_A) * 4 hours distance_B = (rate_B) * 4.5 hours rate_B = rate_A - 5 distance_A = distance_B call rate_A = x x*4 = (x-5)*4.5 4x = 4.5x - 5*4.5 4x = 4.5x - 22.5 .5x = 22.5 x = 22.5 / .5 = 45

OpenStudy (perl):

can you solve it now?

OpenStudy (anonymous):

ok ya that helped me alot thanks!!!!

OpenStudy (perl):

look this over, take your time